# June 2, 2025

### v1

@@
University of Florida, Gainesville,
Florida, 32611, USA
The limits on lepton radii and contact interaction scales are used to derive limits on the deviations of lepton anomalous magnetic dipole moments (g-2)/2 from the Standard Model predictions. In the case of quadratic dependence of the deviations on the lepton mass the limits for electrons are better compared to low energy measurements, for muons somewhat weaker than the latest precision experiments, and for taus by far the best.
@@
At the energy frontier, the lepton-pair measurements at LEP at energies above the Z resonance mass provide very sensitive probes of the point-like structure of the electron [1, 2], the muon and the tau [3].
At the precision frontier, the Fermilab Muon g-2 experiment [4] has generated excitement when comparing their measurements to the best available theory predictions. If the anomalous magnetic moments deviate from theory, this may be a sign for lepton substructure and finite lepton radii, or for new interactions. In this study we analyze the interplay between lepton radii, contact interactions and anomalous magnetic moments.
The high precision measurements of the magnetic dipole moment $\rm(g-2)_{e}$ of the electron have been used to put stringent limits on the electron radius $\rm r_{e}$ [5, 6]. If non-standard contributions to $\rm(g-2)_{e}$ scale linearly with the electron mass, the bound is of the order of $\rm r_{e}\sim 10^{-23}\ m$. As a natural consequence of chiral symmetry, these corrections are expected to scale quadratically with the lepton mass [5]. In this case the bound is reduced to $\rm r_{e}\sim 10^{-18}\ m$.
@@
This limit is substantially lower than the limit derived from $\rm(g-2)_{e}$ measurements in the case where the deviations from the SM of the magnetic dipole moment of the electron depend quadratically on its mass.
@@
As the LEP experiments have observed no statistically significant deviations from the SM, they have derived one-sided lower limits on the scale $\rm\Lambda$ of contact interactions at 95% confidence level [8] under the assumption of a strong coupling $\rm\frac{g^{2}}{4\pi}=1$. The best limits are summarized in Table 2.
@@
In contrast, for contact interactions very stringent limits $\sim$ 15-20 TeV were derived. By convention, a strong coupling $\rm\frac{g^{2}}{4\pi}=1$ for the novel interactions is postulated, which may be too optimistic. In a more conservative approach, if we assume an electromagnetic strength coupling, the limits are reduced substantially:
@@
where $\alpha_{QED}$ at the Z resonance is used and small effects, at the percent level, of the running of $\rm\alpha_{QED}$ to higher energies are ignored. $\lambda$ is the “QED modified”, lower, scale. A careful study shows that the VV contact interaction model with positive interference modifies the differential cross section in a similar way as a form factor, so we will use these results for comparisons. The results for $\Lambda$ and $\lambda$ are compiled in Table 2. As can be seen, in the case of $\lambda$ the bounds on the lepton radii are close to the form factor bounds.
@@
The situation is reversed for tau leptons: as experimental measurements of $\delta a_{\tau}$ are very difficult, our limits by far exceed direct approaches.
@@
To conclude: we used the limits on lepton radii and contact interaction scales to derive limits on the deviations of lepton anomalous magnetic dipole moments (g-2)/2 from the Standard Model predictions. In the case of quadratic dependence of the deviations on the lepton mass the limits for electrons are better compared to low energy measurements, for muons somewhat weaker than the latest precision experiments, and for taus by far the best.

### v2

@@
University of Florida
Gainesville, Florida, 32611, USA
The limits on lepton radii and contact interaction scales are used to derive limits on the deviations of lepton anomalous magnetic dipole moments a=(g-2)/2 from the Standard Model predictions. In the case of quadratic dependence of the deviations on the lepton mass, the limits for electrons are better compared to low energy measurements, for muons somewhat weaker than the latest precision experiments, and for taus by far the best.
@@
At the energy frontier, the lepton-pair measurements at LEP at energies above the Z resonance mass provide very sensitive probes of the point-like structure of the electron [1, 2], the muon, and the tau [3].
At the precision frontier, the Fermilab Muon g-2 experiment [4, 5] has generated excitement when comparing their measurements to the best available theory predictions. If the anomalous magnetic moments deviate from theory, this may be a sign for lepton substructure and finite lepton radii, or for new interactions. In this study we analyze the interplay between lepton radii, contact interactions and anomalous magnetic moments.
The high precision measurements of the magnetic dipole moment $\rm g_{e}$ of the electron have been used to put stringent limits on the electron radius $\rm r_{e}$ [6, 7]. If non-standard contributions to $\rm g_{e}$ scale linearly with the electron mass, the most precise bound [8] is below $\rm 10^{-24}\ m$. As a natural consequence of chiral symmetry, these corrections are expected to scale quadratically with the lepton mass [6]. In this case the bound is reduced to $\rm r_{e}<3.2\cdot 10^{-19}\ m$.
@@
This limit is below the limit derived from $\rm(g-2)_{e}$ measurements in the case where the deviations from the SM of the magnetic dipole moment of the electron depend quadratically on its mass.
@@
As the LEP experiments have observed no statistically significant deviations from the SM, they have derived one-sided lower limits on the scale $\rm\Lambda$ of contact interactions at 95% confidence level [10] under the assumption of a strong coupling $\frac{g^{2}}{\rm{4\pi}}=1$. The best limits are summarized in Table 2.
@@
In contrast, for contact interactions very stringent limits $\sim$ 15-20 TeV were derived. By convention, a strong coupling $\frac{g^{2}}{\rm{4\pi}}=1$ for the novel interactions is postulated, which may be too optimistic. In a more conservative approach, if we assume a coupling of electromagnetic strength, the limits are reduced substantially:
@@
where the value of $\rm\alpha_{QED}$ at the Z resonance is used and small effects, at the percent level, of the running of $\rm\alpha_{QED}$ to higher energies are ignored. $\lambda$ is the “QED modified”, lower, scale. A careful study shows that the VV contact interaction model with positive interference modifies the differential cross section in a similar way as a form factor, so we will use these results for comparisons. The results for $\Lambda$ and $\lambda$ are compiled in Table 2. As can be seen, in the case of $\lambda$ the bounds on the lepton radii are close, and lower, than the form factor bounds.
@@
The situation is reversed for tau leptons: as experimental measurements [12] of $\delta a_{\tau}$ are very difficult, our limits by far exceed direct approaches.
@@
To conclude: we used the limits on lepton radii and contact interaction scales to derive limits on the deviations of lepton anomalous magnetic dipole moments (g-2)/2 from the Standard Model predictions. If the deviations exhibit a quadratic dependence on the lepton mass, the limits for electrons are better compared to low energy measurements, for muons somewhat weaker than the latest precision experiments, and for taus by far the best.